# Multilevel lattice codes from Hurwitz quaternion integers

### v2

@@ IV Multilevel Decoding Process
Let $\mathbf{y}\in\mathbb{R}^{4n}$ be the received point, $\mathbf{y}=\mathbf{x}+\mathbf{w},$ where $\mathbf{x}\in\Lambda_{\pi_{A}}^{\mathscr{H}}(\mathscr{C})$ and $\mathbf{w}\in\mathbb{R}^{4n}$ represents the noise. Since, $\mathbf{x}$ belongs to Construction $\pi_{A}$ lattice, it can be decomposed as
$\mathbf{x}$ $=\left[\sum_{i=1}^{k}\nu_{i}\zeta_{i}\gamma_{i}(\overline{\pi}_{i}a_{i}^{(1)}+\pi_{i}a_{i}^{(2)})\right]\mod q\mathscr{H}+q\mathbf{\tilde{z}},$ (10)
where $q=\prod_{l=1}^{k}p_{l}$.
Consider =$\mathscr{B}=\{v_{1},...,v_{4n}\}$ a basis of $\mathscr{H}^{n}$ and the fundamental region $\mathscr{P}(\mathscr{B})$ of $q\mathscr{H}^{n}$ in $\mathbb{R}^{4n}$, $\mathscr{P}(\mathscr{B})=\{\sum_{i=1}^{4n}a_{i}v_{i},a_{i}\in[-q/2,q/2)\},$ then reducing $\mathbf{y}$ to $\mathscr{P}(\mathscr{B})$, denoted by $\mathbf{y}\mod q\mathscr{H}$, we get,
$\mathbf{y}\mod q\mathscr{H}=\mathbf{y}-q\hat{z}=y_{1}^{(1)}.$
where $\hat{z}=\lfloor q^{-1}\cdot y\rceil$.
@@
$y_{1}^{(1)}=\left[\sum_{i=1}^{k}\nu_{i}\zeta_{i}\gamma_{i}(\overline{\pi}_{i}a_{i}^{(1)}+\pi_{i}a_{i}^{(2)})\right]\mod q\mathscr{H}+q\mathbf{z}+\mathbf{w},$
where $\mathbf{z}=\mathbf{\tilde{z}}-\mathbf{\hat{z}}\in\mathscr{H}^{n}$.
@@
$y_{1}^{\prime(1)}=$ $y_{1}^{(1)}\mod\pi_{1}\mathscr{H}=\nu_{1}\zeta_{1}\gamma_{1}\overline{\pi}_{1}a_{1}^{(1)}\mod\pi_{1}\mathscr{H}+\mathbf{w}_{1}$
$=$ $a_{1}^{(1)}\mod\pi_{1}\mathscr{H}+\mathbf{w}_{1}^{(1)},$
where $\mathbf{w}_{1}^{(1)}=\mathbf{w}\mod\pi_{1}\mathscr{H}$.
Apply a decoder to $y_{1}^{\prime(1)}$ to obtain the codeword $\mathbf{a_{1}^{(1)}}\in\mathscr{C}_{1}^{(1)}$. Once we have $\mathbf{a_{1}^{(1)}}$, perform a re-encoding process by
$y_{1}^{(2)}=y_{1}^{(1)}-\nu_{1}\zeta_{1}\gamma_{1}\overline{\pi}_{1}\mathbf{a_{1}^{(1)}}.$
Repeat the same steps, now operating with $\mathscr{C}_{1}^{(2)}$, applying $\mod\overline{\pi}_{1}\mathscr{H}$ to $y_{1}^{(2)}$ to get $\mathbf{a_{1}^{(2)}}$ and then re-encoding again to get
$y_{2}^{(1)}=y_{1}^{(2)}-\nu_{1}\zeta_{1}\gamma_{1}\pi_{1}\mathbf{a_{1}^{(2)}}.$
Repeat these steps for all levels $L=2,\ldots,k$.
In the last step $k+1$, we have
$y_{k+1}=\mathbf{w}+q\mathbf{z}\Rightarrow\mathbf{w}=y_{k+1}\mod q\mathscr{H},$
Then we can recover $\mathbf{x}$, the lattice point closest to the received $\mathbf{y}$,
$\mathbf{x}=\mathbf{y}-\mathbf{w}.$
Algorithm 1 is the description of this process. We denote as $\text{Dec}_{j},j=1,...,k$, any decoder that identifies the closest point $\mathbf{x}$ in the lattice from $\mathscr{C}_{i}^{(j)}$.
@@
0:  received message $\mathbf{y}$.
  $y_{0}=\mathbf{y}\mod q\mathscr{H}$;
  $y_{0}=y_{1}$;
  for $j=1$ to $k$ do
@@
  $\mathbf{w}=y_{k+1}\mod q\mathscr{H}$;
  $\mathbf{x}=\mathbf{y}-\mathbf{w}$
  return  estimated lattice point $\mathbf{x}\in\Lambda_{\pi_{A}}^{\mathscr{H}}(\mathscr{C})$.
@@ IV-A Decoding complexity
Using Construction $\pi_{A}$ over Hurwitz integers, we can improve the decoding process particularly for codes with “large” sizes. It is illustrated in Table 1 the reduction of the computational time for decoding codes over $(\mathscr{H}/q\mathscr{H})^{2}$ using the serial modulo decoding (SMD) compared to standard maximum likelihood decoding (MLD). The comparison was performed considering Construction $\pi_{A}$ (SMD) and Construction $A$ (MLD) from the same code. The experiments were performed using an Intel® Xeon® Gold 6154 CPU and was implemented using Wolfram Mathematica [43].
TABLE I: Time comparison using decoding algorithm MLD and SMD for codes over $(\mathscr{H}/q\mathscr{H})^{2}$, the time was measured in seconds. All codes have $\text{rank}(\mathscr{C})=1$ and size $|\mathscr{C}|=q^{4}$.
$q$ Code Size MLD SMD
3 81 0.47028 0.003784
5 625 26.2844 0.005913
7 2401 407.118 0.006782
11 14641 15656.1 0.011025
13 28561 60254.0 0.011632
15 50625 199020.0 0.010448
17 83521 547500.0 0.028623
The maximum-likelihood decoder (MLD) over an AWGN channel works by finding the nearest neighbor to the received word. This essentially involves obtaining the minimal difference vector between the received word and the codewords of the codebook. For coding over $\mathbb{F}_{q}$, as in [1], we assume that a $q-$ary LDPC code is implemented for which the decoding complexity is reported to be roughly $O(q\log_{2}(q))$, with $q=p^{b}$ a power of a prime number [44].
In the context of a multilevel decoder, maximum-likelihood decoding occurs independently at each level. Consequently, the decoding process is dominated by the largest cardinality of the codes involved in the encoding.
@@
$(\mathscr{R}/q\mathscr{R})^{n}\cong(\mathscr{R}/p_{1}\mathscr{R})^{n}\times\dots(\mathscr{R}/p_{r}\mathscr{R})^{n}$
with $q=p_{1}\cdot\dots\cdot p_{r}$, where $p_{i}$ are prime elements over $\mathscr{R}=\mathbb{Z},\mathbb{Z}[i],\mathbb{Z}[\omega]$, or $\mathscr{H}$, we have
$\mathscr{C}=\varphi(\mathscr{C}_{1}\times...\times\mathscr{C}_{r})\subset(\mathscr{R}/q\mathscr{R})^{n}$
where $\mathscr{C}_{i}\subset(\mathscr{R}/p_{i}\mathscr{R})^{n}$ represents a linear code of size $|\mathscr{C}_{i}|$, and $|\mathscr{C}|=\prod_{i=1}^{r}|\mathscr{C}_{i}|$. Thus, $\mathscr{C}$ is associated with a lattice code in $\mathbb{R}^{n},\mathbb{R}^{2n},\mathbb{R}^{2n}$, and $\mathbb{R}^{4n}$ for $\mathscr{R}=\mathbb{Z},\mathbb{Z}[i],\mathbb{Z}[\omega]$, and $\mathscr{H}$, respectively. The decoding complexity of Construction $\pi_{A}$ is therefore $O(|\mathscr{C}_{\text{max}}|\log_{2}|\mathscr{C}_{\text{max}}|)$ dominated by $|\mathscr{C}_{\text{max}}|$, where $|\mathscr{C}_{\text{max}}|=\max\{|\mathscr{C}_{i}|\}$ [1], [15].
It’s worth noting that using lattices constructed by Hurwitz integers also expands the design space. For instance, for a code $\mathscr{C}$ with $|\mathscr{C}|=81$, there is a Construction $\pi_{A}$ over the Hurwitz integers as seen in Example 4, but there is no Construction $\pi_{A}$ over $\mathbb{Z},\mathbb{Z}[i]$, or $\mathbb{Z}[\omega]$ due to the lack of a partition into distinct prime ideals for a code of this cardinality. Moreover, also for $p\equiv 11\mod 12$, a partition always exists for Hurwitz integers but codes with cardinality $p^{4}$ cannot be obtained by Construction $\pi_{A}$ over $\mathbb{Z},\mathbb{Z}[i]$ or $\mathbb{Z}[\omega]$.
Moreover, considering codes $\mathscr{C}\in(\mathscr{R}/q\mathscr{R})^{n}$ with $|\mathscr{C}|=|(\mathscr{R}/q\mathscr{R})^{n}|$ obtained through Construction $\pi_{A}$ in dimension $4n$, the computational complexity of decoding the lattice codes over $\mathbb{Z},\mathbb{Z}[i]$, and $\mathbb{Z}[\omega]$ do not exceed those obtained through $\mathscr{H}$ as detailed next.
@@
$\left(\frac{\mathbb{Z}}{q\mathbb{Z}}\right)^{4n}\cong\left(\frac{\mathbb{Z}}{p_{1}\mathbb{Z}}\right)^{4n}\times\left(\frac{\mathbb{Z}}{p_{2}\mathbb{Z}}\right)^{4n}\times\ldots\times\left(\frac{\mathbb{Z}}{p_{r}\mathbb{Z}}\right)^{4n},$
and the computational complexity is $O(p_{\mathbb{Z}}\log_{2}p_{\mathbb{Z}})$, where $p_{\mathbb{Z}}=\max\{p_{i}^{4n}\}$, $i=1,...,r$.
For the Gaussian integers, $\mathscr{R}=\mathbb{Z}[i]$, we can rearrange $q=p_{1}\cdot\ldots\cdots p_{s}\cdot p_{s+1}\cdot\ldots\cdot p_{r}$, with $p_{i}\equiv 1\mod 4,i=1,...,s$ and $p_{j}\equiv 3\mod 4,j=s+1,...,r$. Since $\mathbb{Z}[i]/p\mathbb{Z}[i]$ is factorable only for $i=1,...,s$, i.e, $(\mathbb{Z}[i]/p_{i}\mathbb{Z}[i])^{2n}=(\mathbb{Z}[i]/\pi_{i}\mathbb{Z}[i])^{2n}\times(\mathbb{Z}[i]/\overline{\pi}_{i}\mathbb{Z}[i])^{2n}$, we have
@@
×$\times\left(\frac{\mathbb{Z}[i]}{p_{s+1}\mathbb{Z}[i]}\right)^{2n}\times\ldots\times\left(\frac{\mathbb{Z}[i]}{p_{r}\mathbb{Z}[i]}\right)^{2n},$
the computational complexity is $O(p_{G}\log_{2}p_{G})$, where $p_{G}=\max\{p_{i}^{2n},p_{j}^{4n}\}$, $i=1,...,s$ and $j=s+1,...,r$. We can proceed in an analogous way in the case of Eisenstein integers, rearrange $q=p_{1}\cdot\ldots\cdots p_{\tilde{s}}\cdot p_{\tilde{s}+1}\cdot\ldots\cdot p_{r}$, with $p_{i}\equiv 1\mod 3,i=1,...,\tilde{s}$ and $p_{j}\equiv 2\mod 3,j=\tilde{s}+1,...,r$. Again, we have that $(\mathbb{Z}[\omega]/p_{i}\mathbb{Z}[\omega])^{2n}$ is only factorable for $i=1,...,\tilde{s}$ and we can assure that the computational complexity is $O(p_{E}\log_{2}p_{E})$, where $p_{E}=\max\{p_{i}^{2n},p_{j}^{4n}\}$, $i=1,...,\tilde{s}$ and $j=\tilde{s}+1,...,r$.
@@
with computational complexity $O(p_{\mathscr{H}}\log_{2}p_{\mathscr{H}})$, where $p_{\mathscr{H}}=\max\{p_{i}^{2n}\}$, $i=1,...,r$.
@@
Table II and Figure 1 compare the computational complexity of decoding in Construction $\pi_{A}$ over $\mathbb{Z}[i]$, $\mathbb{Z}[\omega]$, and $\mathscr{H}$, in real dimension $4$ (i.e. $n=1$) and constellations with cardinality $q^{4}$.
@@
$q$
Factorization
3
$\mathscr{H}$
$(1+i+j)\cdot(1-i-j)$

### v3

@@ IV Multilevel Decoding Process
Let $y\in\mathbb{R}^{4n}$ be the received point,
$y=x+z,$
where $x\in\Lambda_{\pi_{A}}^{\mathscr{H}}(\mathscr{C})$, and $z\in\mathbb{R}^{4n}$ represents the additive white Gaussian noise with variance $\sigma^{2}$. Since $x$ belongs to Construction $\pi_{A}$ lattice, it can be decomposed as
$x$ $=\left[\sum_{i=1}^{r}\nu_{i}\zeta_{i}\gamma_{i}(\overline{\pi}_{i}a_{i}^{(1)}+\pi_{i}a_{i}^{(2)})\mod q\mathscr{H}\right]+q\tilde{\lambda},$ (10)
where $q=\prod_{l=1}^{r}p_{l}$ and $\tilde{\lambda}\in\mathscr{H}^{n}$.
Consider =$\mathscr{B}=\{v_{1},\ldots,v_{4n}\}$ a basis of $\mathscr{H}^{n}$ and the fundamental region $\mathscr{P}(\mathscr{B})$ of $q\mathscr{H}^{n}$ in $\mathbb{R}^{4n}$, $\mathscr{P}(\mathscr{B})=\{\sum_{i=1}^{4n}a_{i}v_{i},a_{i}\in[-q/2,q/2)\},$ then reducing $y$ to $\mathscr{P}(\mathscr{B})$, denoted by $y\mod q\mathscr{H}$, we get
$y\mod q\mathscr{H}=y-q\hat{\lambda}=:y_{1}^{(1)},$
where $\hat{\lambda}=\lfloor q^{-1}\cdot y\rceil$.
@@
$y_{1}^{(1)}=\left[\sum_{i=1}^{r}\nu_{i}\zeta_{i}\gamma_{i}(\overline{\pi}_{i}a_{i}^{(1)}+\pi_{i}a_{i}^{(2)})\mod q\mathscr{H}\right]+q\lambda+z,$
where $\lambda:=\tilde{\lambda}-\hat{\lambda}\in\mathscr{H}^{n}$.
@@
$y_{1}^{\prime(1)}:=$ $y_{1}^{(1)}\mod\pi_{1}\mathscr{H}=\nu_{1}\zeta_{1}\gamma_{1}\overline{\pi}_{1}a_{1}^{(1)}\mod\pi_{1}\mathscr{H}+z_{1}$
$=$ $a_{1}^{(1)}\mod\pi_{1}\mathscr{H}+z_{1}^{(1)},$
where $z_{1}^{(1)}:=z\mod\pi_{1}\mathscr{H}$.
Apply a decoder to $y_{1}^{\prime(1)}$ to obtain the codeword $a_{1}^{(1)}\in\mathscr{C}_{1}^{(1)}$. Once we have $a_{1}^{(1)}$, perform a re-encoding process by
$y_{1}^{(2)}:=y_{1}^{(1)}-\nu_{1}\zeta_{1}\gamma_{1}\overline{\pi}_{1}a_{1}^{(1)}.$
Repeat the same steps, now operating with $\mathscr{C}_{1}^{(2)}$, applying $\mod\overline{\pi}_{1}\mathscr{H}$ to $y_{1}^{(2)}$ to get $a_{1}^{(2)}$ and then re-encoding again to get
$y_{2}^{(1)}:=y_{1}^{(2)}-\nu_{1}\zeta_{1}\gamma_{1}\pi_{1}a_{1}^{(2)}.$
Repeat these steps for all levels $L=2,\ldots,r$.
In the last step $r+1$, we have
$y_{r+1}=z+q\lambda\Rightarrow z=y_{r+1}\mod q\mathscr{H},$
Then we can recover $x$, the lattice point closest to the received $y$,
$x=y-z.$
Algorithm 1 is the description of this process. We denote as $\text{Dec}_{j},j=1,\ldots,r$, any decoder that identifies the closest point $x$ in the Construction A lattice from $\mathscr{C}_{i}^{(j)}$.
@@
0:  received message $y$.
  $y_{1}=y\mod q\mathscr{H}$;
  $y_{1}^{(1)}=y_{1}$;
  for $j=1$ to $r$ do
@@
  $z=y_{r+1}\mod q\mathscr{H}$;
  $x=y-z$
  return  estimated lattice point $x\in\Lambda_{\pi_{A}}^{\mathscr{H}}(\mathscr{C})$.
@@ IV-A Decoding complexity
Using Construction $\pi_{A}$ over Hurwitz integers, we can enhance the decoding complexity for codes of “large” sizes. It is noteworthy that, related to Construction $\pi_{A}$ lattices over $\mathbb{Z}$, $\mathbb{Z}[\mathrm{i}]$, $\mathbb{Z}[\omega]$, [3], the Construction $\pi_{A}$ over $\mathscr{H}$ for $\mathscr{C}=\varphi^{-1}(\mathscr{C}_{1}^{(1)}\times\ldots\times\mathscr{C}_{r}^{(2)})$ has decoding complexity influenced by the greatest cardinality of codes $\mathscr{C}_{1}^{(1)},\mathscr{C}_{1}^{(2)},\ldots,\mathscr{C}_{r}^{(1)},\mathscr{C}_{r}^{(2)}$, denoted as $M=\max\{|\mathscr{C}_{i}^{(1)}|,|\mathscr{C}_{i}^{(2)}|,i=1,\ldots,r\}$, roughly $O(M\log_{2}M)$.
Table I presents a comparison of the computational time required to decode linear codes generated through Construction $\pi_{A}$ over $(\mathbb{Z}/q\mathbb{Z})^{4}$, $(\mathbb{Z}[\mathrm{i}]/q\mathbb{Z}[\mathrm{i}])^{2}$, $(\mathbb{Z}[\omega]/q\mathbb{Z}[\omega])^{2}$, and $\mathscr{H}/q\mathscr{H}$. The decoding was performed using the serial modulo decoding algorithm. For a fair comparison, all codes considered have the same size, $q^{4}$ elements in the respective quotient ring, and we used the canonical basis vectors as generators in each case. A notable improvement in computational efficiency is observed in some cases when the decoding over $\mathscr{H}$ is lower compared to the other rings. However, it is worth mentioning that for $q=39$, where the maximum code size in both $\mathbb{Z}[i]/q\mathbb{Z}[i]$ and $\mathscr{H}/q\mathscr{H}$ is $13^{2}$, the decoding over $\mathbb{Z}[i]$ requires only $0.54506$ seconds, while decoding over $\mathscr{H}$ takes $1.22351$ seconds. All simulations were performed using an Intel® Xeon® Gold 6154 CPU.
TABLE I: Decoding time comparison between Construction $\pi_{A}$ lattices from $\mathbb{Z}$, $\mathbb{Z}[\mathrm{i}]$, $\mathbb{Z}[\omega]$, and $\mathscr{H}$.
Time (seconds)
[…]
15 $50625$ $3.50723$ $0.080691$ 0.060011
21 $194481$ $62.5559$ $0.53549$ 0.234559
33 $1185921$ $2600.66$ 0.681975
35 $1500625$ $66.3755$ $100.837$ $21.7788$ 0.663997
39 $2313441$ $8718.13$ 0.54506 $1.22351$
55 $9150625$ $2328.67$ $4014.38$ 0.684826
77 $35153041$ $2576.51$ $15284.9$ 0.762427
As remarked in [3], for Construction A lattices the decoding complexity is dominated by $|\mathscr{C}|$. In the context of a Construction $\pi_{A}$ multilevel decoder, maximum-likelihood decoding occurs independently at each level. Consequently, the decoding process is dominated by the largest cardinality of the codes involved in the encoding.
@@
$(\mathscr{R}/q\mathscr{R})^{n}\cong(\mathscr{R}/p_{1}\mathscr{R})^{n}\times\cdots\times(\mathscr{R}/p_{r}\mathscr{R})^{n},$
with $q=p_{1}\cdot\ldots\cdot p_{r}$, where $p_{i}$ are prime elements over $\mathscr{R}=\mathbb{Z},\mathbb{Z}[i],\mathbb{Z}[\omega]$, or $\mathscr{H}$, we have
$\mathscr{C}=\varphi^{-1}(\mathscr{C}_{1}\times\cdots\times\mathscr{C}_{r})\subset(\mathscr{R}/q\mathscr{R})^{n},$
where $\mathscr{C}_{i}\subset(\mathscr{R}/p_{i}\mathscr{R})^{n}$ represents a linear code of size $|\mathscr{C}_{i}|$, and $|\mathscr{C}|=\prod_{i=1}^{r}|\mathscr{C}_{i}|$. Thus, $\mathscr{C}$ is associated with a lattice code in $\mathbb{R}^{n},\mathbb{R}^{2n},\mathbb{R}^{2n}$, and $\mathbb{R}^{4n}$ for $\mathscr{R}=\mathbb{Z},\mathbb{Z}[i],\mathbb{Z}[\omega]$, and $\mathscr{H}$, respectively. For LDPC codes in each level, the decoding complexity of Construction $\pi_{A}$ is reported to be roughly $O(\mathscr{C}_{\text{max}}\log_{2}\mathscr{C}_{\text{max}})$, where $\mathscr{C}_{\text{max}}=\max\{|\mathscr{C}_{i}|\}$, [3], [44].
It’s worth noting that using lattices constructed from Hurwitz integers also expands the design space. For instance, for a code $\mathscr{C}$ with $|\mathscr{C}|=81$, there is a Construction $\pi_{A}$ over the Hurwitz integers, but there is no Construction $\pi_{A}$ over $\mathbb{Z},\mathbb{Z}[i]$, or $\mathbb{Z}[\omega]$ due to the lack of a partition into distinct prime ideals for a code of this cardinality. Moreover, also for $p\equiv 11\mod 12$, a partition always exists for Hurwitz integers but codes with cardinality $p^{4}$ cannot be obtained by Construction $\pi_{A}$ over $\mathbb{Z},\mathbb{Z}[i]$ or $\mathbb{Z}[\omega]$.
Additionally, considering codes $\mathscr{C}=(\mathscr{R}/q\mathscr{R})^{4n}$ obtained through Construction $\pi_{A}$ in dimension $4n$, the computational complexity of decoding the lattice codes over $\mathbb{Z},\mathbb{Z}[i]$, and $\mathbb{Z}[\omega]$ do not exceed those obtained through $\mathscr{H}$ as detailed next.
@@
$\left(\frac{\mathbb{Z}}{q\mathbb{Z}}\right)^{4n}\cong\left(\frac{\mathbb{Z}}{p_{1}\mathbb{Z}}\right)^{4n}\times\left(\frac{\mathbb{Z}}{p_{2}\mathbb{Z}}\right)^{4n}\times\cdots\times\left(\frac{\mathbb{Z}}{p_{r}\mathbb{Z}}\right)^{4n},$
and the computational complexity is $O(p_{\mathbb{Z}}\log_{2}p_{\mathbb{Z}})$, where $p_{\mathbb{Z}}=\max\{p_{i}^{4n}\}$, $i=1,\ldots,r$.
For the Gaussian integers, $\mathscr{R}=\mathbb{Z}[i]$, we can rearrange $q=p_{1}\cdot\ldots\cdot p_{s}\cdot p_{s+1}\cdot\ldots\cdot p_{r}$, with $p_{i}\equiv 1\mod 4,i=1,\ldots,s$ and $p_{j}\equiv 3\mod 4,j=s+1,\ldots,r$. Since $\mathbb{Z}[i]/p_{i}\mathbb{Z}[i]$ is factorable in non trivial rings only for $i=1,\ldots,s$, i.e, $(\mathbb{Z}[i]/p_{i}\mathbb{Z}[i])^{2n}=(\mathbb{Z}[i]/\pi_{i}\mathbb{Z}[i])^{2n}\times(\mathbb{Z}[i]/\overline{\pi}_{i}\mathbb{Z}[i])^{2n}$, we have
@@
×$\times\left(\frac{\mathbb{Z}[i]}{p_{s+1}\mathbb{Z}[i]}\right)^{2n}\times\ldots\times\left(\frac{\mathbb{Z}[i]}{p_{r}\mathbb{Z}[i]}\right)^{2n}$
$\cong$ $\left(\frac{\mathbb{Z}[i]}{\pi_{1}\mathbb{Z}[i]}\right)^{2n}\times\left(\frac{\mathbb{Z}[i]}{\overline{\pi}_{1}\mathbb{Z}[i]}\right)^{2n}\times\ldots\times\left(\frac{\mathbb{Z}[i]}{\pi_{r}\mathbb{Z}[i]}\right)^{2n}\times$
$\left(\frac{\mathbb{Z}[i]}{\overline{\pi}_{r}\mathbb{Z}[i]}\right)^{2n}\times\left(\frac{\mathbb{Z}[i]}{p_{s+1}\mathbb{Z}[i]}\right)^{2n}\times\ldots\times\left(\frac{\mathbb{Z}[i]}{p_{r}\mathbb{Z}[i]}\right)^{2n},$
the computational complexity is $O(p_{G}\log_{2}p_{G})$, where $p_{G}=\max\{p_{i}^{2n},p_{j}^{4n},i=1,\ldots,s,\ j=s+1,\ldots,r\}$. We can proceed in an analogous way in the case of Eisenstein integers, by rearranging $q=p_{1}\cdot\ldots\cdots p_{\tilde{s}}\cdot p_{\tilde{s}+1}\cdot\ldots\cdot p_{r}$, with $p_{i}\equiv 1\mod 3,i=1,\ldots,\tilde{s}$ and $p_{j}\equiv 2\mod 3,j=\tilde{s}+1,\ldots,r$. Again, we have that $(\mathbb{Z}[\omega]/p_{i}\mathbb{Z}[\omega])^{2n}$ is only factorable on non trivial rings for $i=1,\ldots,\tilde{s}$, and we can assure that the computational complexity is $O(p_{E}\log_{2}p_{E})$, where $p_{E}=\max\{p_{i}^{2n},p_{j}^{4n},i=1,\ldots,\tilde{s},\ j=\tilde{s}+1,\ldots,r\}$.
@@
with computational complexity $O(p_{\mathscr{H}}\log_{2}p_{\mathscr{H}})$, where $p_{\mathscr{H}}=\max\{p_{i}^{2n}\}$, $i=1,\ldots,r$.
@@
Table II and Figure 2 compare the computational complexity of decoding in Construction $\pi_{A}$ over $\mathbb{Z}[i]$, $\mathbb{Z}[\omega]$, and $\mathscr{H}$, in real dimension $4$ (i.e. $n=1$) and for constellations with cardinality $q^{4}$.
Figure 2: Comparison of decoding complexity for constellations with cardinality $q^{4}$ in $\mathbb{R}^{4}$ obtained through Construction $\pi_{A}$ over $\mathbb{Z}[i]$, $\mathbb{Z}[\omega]$, and $\mathscr{H}$. Here, $\mathscr{C}_{\text{max}}$ denotes the maximum cardinality among the layers in the encoding.
@@
$q$ Factorization $\mathscr{C}_{\text{max}}$
3 $\mathscr{H}$ $(1+i+j)\cdot(1-i-j)$ $3^{2}$